COMBINATORIAL REPRESENTATION THEORY

HÉLÈNE BARCELO ARUN RAM

Abstract. We attempt to survey the field of combinatorial representation theory, describe the main results and main questions and give an update of its current status. We give a personal viewpoint on the field, while remaining aware that there is much important and beautiful work that we have not been able to mention.